Grade 12General Physics

kinetic energy of a satellite is x then what is the time period proportional to

SHREYA GAUTAM
10 Years agoGrade 12

1 Answer

Vikas TU
10 Years ago
Kinetic enrergy for satellite is given as:
1/2mv^2 =  x
now, v = wr
where w is the angular speed os satellite.
1/2*m*(wr)^2 = x 
T ime Period, T = 2pi/w or w = 2pi/T
1/2*m*(2pi*r/T)^2 = x
or
T^2 = pi^2*r^2*m/x
or
x = pi^2*r^2*m/T^2
or
T^2 = C/x
T^2 is inversely proprtional to x.